# Giovanni Compiani Ilya Morozov Stephan Seiler

### v3

@@
Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Jon McClure, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments. We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance. Contact information: Giovanni Compiani (gio1compiani@gmail.com), Ilya Morozov (ilya.morozov@kellogg.northwestern.edu), Stephan Seiler (stephan.a.seiler@gmail.com). Giovanni Compiani gratefully acknowledges support from the Center for Applied AI at Booth.
@@
We propose a demand estimation method that leverages unstructured data to infer substitution patterns. Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a mixed logit demand model. This approach enables demand estimation even when researchers lack data on product attributes or when consumers value hard-to-quantify attributes, such as visual design. Using a choice experiment, we show this approach outperforms standard attribute-based models at counterfactual predictions of second choices. We also apply it to 40 product categories on Amazon.com and consistently find that unstructured data identifies close substitutes within each category.
@@ 1 Introduction
Many problems in economics and marketing—such as merger analysis, tariff evaluation, and optimal pricing—require researchers to estimate demand for differentiated products. A standard approach to these problems has been to estimate demand models that capture substitution through the similarity of product attributes. While common, this approach faces two practical challenges. First, researchers rarely observe all attributes that differentiate products. Instead, they rely on third-party data, where attributes are chosen based on unknown criteria, or gather their own data, subjectively selecting which attributes to collect. The collected attributes may not align with those most relevant to consumer choices. Second, consumers often consider visual design and functional benefits of products—dimensions that are difficult to capture through observed attributes. [^1] This criticism of attribute-based models has a long history in the economics literature. For example, Hausman (1994, p.229) skeptically remarks that applying such models to French champagne choices would require researchers to somehow quantify the bubble content.
@@
In this paper, we show how researchers can incorporate text and image data in demand estimation to recover substitution patterns. We propose an approach that uses product images, titles, descriptions, and customer reviews—unstructured data that are widely available even in markets where collecting product attributes is challenging. Using pre-trained deep learning models, we transform images and texts into vector representations—embeddings. We then apply Principal Component Analysis (PCA) to reduce dimensionality and capture the main dimensions of product differentiation. We incorporate these principal components into a mixed logit demand model (Berry et al., 1995; Nevo, 2001), interacting them with random coefficients similar to how researchers usually treat observed product attributes in demand estimation.
@@
This approach enables researchers to incorporate hard-to-quantify product attributes, such as visual design from images and functional benefits from text, while also circumventing the need to subjectively choose which observed attributes to collect. Given the prevalence of unstructured data, we see our approach as a valuable addition to the toolbox of empirical researchers. To make it easier for others to apply our approach, we provide a publicly available Python package, DeepLogit, alongside this paper. [^2] The package is available on PyPI and in the public GitHub repository: github.com/deep-logit-demand/deeplogit. Our package heavily borrows from the existing package Xlogit for GPU-accelerated estimation of mixed logit models, developed by Arteaga et al. (2022).
@@
To show that the proposed method performs well at recovering substitution patterns, we validate it extensively on several datasets. Validation in demand estimation is challenging because substitution patterns are usually unobserved, leaving no clear “ground truth” to evaluate model predictions. To overcome this challenge, we design a choice experiment where participants choose a book from a list of options. We randomize both prices and rankings to identify substitution patterns. Crucially, we elicit both first and second choices, estimate different demand models using only first choices, and evaluate their performance based on how well they predict second choices counterfactually. [^3] We refer to second choice predictions as “counterfactual predictions” in the sense that second choices are not used during estimation, which is based solely on first choices. Conditional second choices are informative about substitution because they reveal which product a consumer would substitute to if their preferred option was unavailable. Recording second choices thus enables us to assess how well each demand model predicts substitution patterns. [^4] Second choices are related to the notion of diversion ratios, which reflect the degree of substitutability and thus the intensity of price competition between two products (Shapiro, 1995; Conlon and Mortimer, 2021; Conlon et al., 2023). Recovering them is crucial for antitrust authorities trying to predict price increases from horizontal mergers and for managers optimizing assortments and prices.
@@
Using these experimental data, we show that both text and images contain useful information about substitution patterns. Intuitively, images capture substitution because book covers visually convey information about genre. Text performs even better, likely because descriptions and reviews capture nuanced details about book plots, helping identify similar books even within the same genre. We then show that the best performing model—based on customer reviews—improves counterfactual predictions of second-choices relative to the standard mixed logit with observed attributes. In fact, the improvement relative to that model is on par with the improvement that the mixed logit achieves over the plain logit model without random coefficients. Thus, in this application, we can recover substitution from unstructured data alone, and in fact can achieve better performance than is possible even with attribute data.
To assess the economic importance of these improvements, we simulate hypothetical mergers of book publishers. We find that differences in estimated substitution patterns are large enough such that an antitrust agency relying on the standard mixed logit with attributes might incorrectly approve mergers that should be challenged and block mergers that should be approved. Thus, our results show that incorporating text and image data yields improvements in estimated substitution patterns that are both statistically and economically significant.
@@
We also validate our approach on observational data. We apply it to 40 product categories on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing. We combine data on online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product detail pages. Here, our contribution is twofold. First, we show that text and images contain useful signals about substitution in all studied categories, highlighting that our approach is broadly applicable across markets. Second, we find—perhaps unexpectedly—that text performs best in some categories where images might be expected to be more useful, such as clothing, and vice versa. A practical takeaway is that researchers may benefit from collecting both text and images, even when only one seems relevant a priori.
@@
Our paper contributes to the extensive literature on demand estimation. Specifically, we build on papers that model substitution through the heterogeneity of preferences over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Nevo, 2001; Berry et al., 2004). We show that incorporating the principal components of text and image embeddings into these models captures substitution patterns well. A key contribution of our paper is that we validate this approach extensively using both experimental and observational data.
@@
This paper is also related to a vast literature in computer science that transforms unstructured data into lower-dimensional representations (Krizhevsky et al., 2012; Mikolov et al., 2013; Goodfellow et al., 2016). There is a growing literature on applications of pre-trained deep learning models in economics and marketing (see Battaglia et al., 2024 for a review). While most prior work focuses on predicting observed outcomes, our goal is fundamentally different: we aim to predict counterfactual quantities, such as substitution patterns. Predicting these quantities well is crucial for applications like merger simulations and optimal pricing. Yet, they are usually unobserved in choice data. As a result, we cannot apply standard cross-validation techniques and instead need to design a custom experiment that allows us to measure these otherwise unobserved quantities. To our knowledge, we are the first to use such an experiment for validation of demand models. [^5] Berry et al. (2004) show that second choice data are useful for estimating substitution patterns, and Conlon et al. (2022) estimate demand from aggregate data on first-choice probabilities and a subset of second-choice probabilities. Conceptually, our validation approach reverses the logic of these papers: we only use first-choice data for estimation and treat second choices as “holdout” data for model validation. Additionally, Raval et al. (2022) use hospital closures induced by natural disasters to compare the ability of various models to predict diversions.
Several other papers have proposed leveraging text and image data in demand estimation. [^6] Netzer et al. (2012) use data on the co-occurrence of products mentioned in online discussion forums to generate a visual representation of competing products, but they do not incorporate these measures into demand estimation. Dew (2024) uses image embeddings from pre-trained models to predict consumers’ ratings for products based on past ratings elicited in a survey. Sisodia et al. (2024) extract interpretable product attributes from images and use them to estimate preferences. Our approach is simpler and scales better across product categories, and thus, it may be preferable in applications where interpretability of extracted attributes is not essential. Quan and Williams (2019) incorporate product image embeddings as shifters of mean utilities. Similarly, Lee (2024) employs large-language models to predict utility intercepts of new products based on their textual descriptions. These methods predict mean utilities but are not designed to estimate utility covariances, which play a crucial role in shaping substitution. [^7] For example, Lee (2024) inverts a plain logit model without random coefficients to extract product fixed effects for the prediction part of his analysis. Thus, his model restricts substitution patterns due to the IIA property. In contrast, our approach models utility covariances and can thus capture flexible substitution patterns—a key input for many empirical applications of interest. Further, while utility intercepts can be recovered from market shares, our goal is to capture substitution patterns that are unobserved. We measure them in an experiment by eliciting second choices, and we use these data for validation. [^8] A recent paper by Han and Lee (2025) uses unstructured data to estimate substitution patterns among fonts in order to study the effects of copyright protection policies.
@@
A possible alternative to our use of unstructured data is to estimate substitution patterns using survey data. For example, Dotson et al. (2019) ask survey participants to rate each product image and then incorporate rating correlations as shifters of utility correlations into a demand model. Similarly, Magnolfi et al. (2022) elicit relative rankings from survey participants (e.g., “Product A is closer to B than to C”), generate embeddings from these rankings, and use them in a random coefficients logit model. Our approach complements these survey-based methods but has the advantage of using only widely available text and image data, thereby avoiding the need for costly category-specific surveys. Another complementary approach is proposed by McClure (2025), who shows how publicly available recommendations data from online platforms can also be used to estimate substitution patterns.
@@
Our approach is well-suited for several empirical applications. Researchers can use it to address a wide range of economics and marketing questions that require accurate and flexible estimates of product substitution. This includes analyzing how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through altering assortments and prices. Additionally, researchers can apply this approach to study how multi-product firms optimize prices and promotions (Hoch et al., 1995; Vilcassim and Chintagunta, 1995). Finally, because our approach circumvents the need to collect distinct attributes for each category, it may be valuable for researchers trying to estimate demand across many categories (Döpper et al., 2024). [^9] Döpper et al. (2024) remark that estimating demand across multiple categories “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach can address this problem if researchers have access to unstructured data.
@@ 2 Proposed Approach
To extract information from product images, we extract their low-dimensional representations—embeddings—using pre-trained deep learning models. We use four convolutional neural networks: VGG19 (Simonyan and Zisserman, 2015), ResNet50 (He et al., 2016), InceptionV3 (Szegedy et al., 2016), and Xception (Chollet, 2017). [^10] These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”). However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings. The key advantage of using pre-trained models is that it reduces the computational burden of estimation while leveraging models trained on large-scale datasets, known for their strong performance in image classification tasks (Russakovsky et al., 2015). Because these models perform well at distinguishing visually similar objects, we expect embeddings to capture the key visual features that differentiate products. Rather than committing to a single model, we perform model selection to determine which one best captures substitution patterns, as described below.
We also extract information from texts, including product titles, descriptions, and customer reviews. We use a bag-of-words model, which represents text as fixed-length vectors based on word counts, as well as a variation with a TF-IDF vectorizer. [^11] TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words. This makes it more likely that text embeddings capture distinctive words that differentiate products from one another. While these count models are relatively simple, we view them as useful benchmarks because they can still detect attributes mentioned in titles, descriptions, and reviews. In addition, we employ two pre-trained deep learning models: Universal Sentence Encoder (USE) (Cer et al., 2018), and BERT Sentence Transformer (ST) (Reimers and Gurevych, 2019). [^12] The model trained by Reimers and Gurevych (2019) is a more efficient version of the widely used BERT network (Devlin et al., 2018). Both USE and ST produce semantically meaningful sentence embeddings and achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017). As a result, they can identify when sellers or consumers describe the same product attributes and functional benefits using similar language, even if the exact words differ.
@@
Although embeddings compress texts and images into lower-dimensional representations, they remain high-dimensional compared to the number of attributes typically used in demand models. For example, incorporating 512-dimensional VGG19 embeddings into a random coefficients logit model would require prohibitively costly numerical integration over all dimensions. We therefore apply PCA to further reduce dimensionality (Backus et al., 2021).
The dimensionality reduction via PCA is attractive for multiple reasons. Raw embeddings are trained on general-purpose datasets to classify images into broad categories, such as “tablet” or “laptop.” However, to estimate demand, we need to analyze consumer choices within a given category. PCA helps us filter out the variation in embeddings necessary for sorting products into categories, allowing us to focus on dimensions most relevant for analyzing substitution within a category. Additionally, principal components are appealing because they are orthogonal to each other, which avoids multicollinearity issues and simplifies estimation of random coefficients in demand models.
@@
We include principal components into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each included component. Since each deep learning model and data type produces different principal components, a key question is which to include in the demand model. [^13] While we could select the specification that best matches observed second choices in our experiment, this strategy is unavailable to most researchers who usually only have first-choice data. We thus propose a model selection algorithm that relies solely on first choices. We perform model selection using in-sample $AIC$. An advantage of using $AIC$ is that it can be interpreted as a bias-corrected estimator of the expected relative Kullback-Leibler (KL) information based on the maximized log-likelihood function (Akaike, 1998). Further, in our experimental results, in-sample $AIC$ strongly correlates with counterfactual performance, justifying its use for model selection (see Section 3.3 where we also implement $BIC$ and cross-validation).
Our model selection algorithm considers a pool of candidate variables consisting of price and the first $P$ principal components. In our application, we choose $P=6$, which collectively account for 70-80% of variance in embeddings (see Appendix Figure A4). We first estimate all possible models with a random coefficient on a single variable and retain the best-fitting model. We then estimate all possible models with two random coefficients and assess whether fit improves relative to the best single-variable model. We continue to increase the number of variables $K$ that have random coefficients until there is no further $AIC$ improvement. We repeat this across all specifications—defined as a combination of a text or image model and data type (e.g., USE Reviews)—and choose the one with the lowest overall $AIC$. The full procedure is given in Algorithm 1.
This algorithm avoids costly combinatorial search over all possible specifications. At the same time, it ensures that we do not rely too heavily on the ordering of principal components, which reflects the explained variance of embeddings but need not be predictive of preferences. For example, the approach might select a model that has random coefficients on the second and fourth principal components, but not the first or third.
@@
9:     Find subset $R_{K}^{*}\in\mathcal{R}$ that minimizes $AIC$ at $AIC_{K}^{*}$
@@
11:       $\text{BestSpec}^{(m)}\leftarrow R_{K}^{*}$ update the best specification
12:       $\text{BestAIC}^{(m)}\leftarrow AIC_{K}^{*}$ update the lowest AICAIC
13:   end while
14: end for
15: Choose the best-fitting specification $m^{*}=\arg\min_{m}\text{BestAIC}^{(m)}$
@@
There are several reasons why texts and images may predict substitution. Consumers may choose based on how a product looks and how it is described by sellers or other consumers. Alternatively, even when consumers do not directly consider visual or textual descriptions, these descriptions may reflect attributes that drive substitution. Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly. Similarly, photos may showcase design features like color and casing style. Our approach extracts this information and uses it to capture substitution.
Given this interpretation, embeddings correlate with substitution patterns because they capture choice-relevant attributes, but the relationship need not be causal: changing product images or descriptions will not necessarily alter substitution patterns. Consequently, we cannot study questions like optimal product design or positioning unless we further unpack the link between embeddings and substitution, which is outside the scope of this paper. By contrast, our method is well-suited for counterfactuals where embeddings can be held fixed, such as optimal pricing or merger simulations. We provide a more detailed discussion of the counterfactuals enabled by our approach in Section 5.2.

### v4

@@
Demand Estimation with Text and Image Datathanks: We thank Steve Berry, Alex Burnap, Matt Gentzkow, Joonhwi Joo, Philip Haile, Yufeng Huang, Jon McClure, Carl Mela, Sanjog Misra, Olivia Natan, Matt Osborne, Katja Seim, Adam Smith, Yuyan Wang, Kevin Williams and seminar participants at Bocconi University, Chicago Booth, London Business School, Stockholm School of Economics, University of Cologne, Nova SBE, Marketing Science Conference 2022, QME Conference 2025, SICS Conference 2024, AI/ML Yale Conference, and SED Conference for helpful comments. We also thank Celina Park, Tanner Parsons, James Ryan, Janani Sekar, Andrew Sharng, Franklin She, Adam Sy, and Vitalii Tubdenov for outstanding research assistance. Contact information: Giovanni Compiani (gio1compiani@gmail.com), Ilya Morozov (ilya.morozov@kellogg.northwestern.edu), Stephan Seiler (stephan.a.seiler@gmail.com). Giovanni Compiani gratefully acknowledges support from the Center for Applied AI at Booth.
@@
We propose a demand estimation approach that leverages unstructured data to infer substitution patterns. Using pre-trained deep learning models, we extract embeddings from product images and textual descriptions and incorporate them into a mixed logit demand model. This approach enables demand estimation even when researchers lack data on product attributes or when consumers value hard-to-quantify attributes such as visual design. Using a choice experiment, we show this approach substantially outperforms standard attribute-based models at counterfactual predictions of second choices. We also apply it to 40 product categories offered on Amazon.com and consistently find that unstructured data are informative about substitution patterns.
@@ 1 Introduction
Many problems in economics and marketing—such as merger analysis, tariff evaluation, and optimal pricing—require estimating demand for differentiated products. A standard approach has been to estimate demand models that capture substitution through the similarity of product attributes. While common, this approach faces two practical challenges. First, researchers rarely observe all choice-relevant attributes. Instead, they rely on third-party data in which attributes are chosen based on unknown criteria, or gather their own data, subjectively choosing which attributes to collect. Second, consumers often consider visual design and functional benefits of products, dimensions that are difficult to capture through observed attributes. [^1] This criticism of attribute-based models has a long history in the economics literature. For example, Hausman (1994, p.229) skeptically remarks that applying such models to French champagne choices would require researchers to somehow quantify the bubble content.
@@
In this paper, we show how researchers can incorporate unstructured data—such as product images, descriptions, or reviews—in demand estimation to recover substitution patterns. Using pre-trained deep learning models, we transform images and texts into vector representations (embeddings) and apply Principal Component Analysis (PCA) to reduce dimensionality and capture the main dimensions of product differentiation. We incorporate these principal components into a logit model with random coefficients, treating them as researchers usually treat observed product attributes. Lastly, we implement a model selection algorithm to choose the type of unstructured data and deep learning model that extract the strongest signal of substitution.
@@
This approach lets researchers incorporate hard-to-quantify product attributes, such as visual design in images and functional benefits from text, while avoiding subjective choices about which attributes to collect. It relies on product images and textual descriptions, which are widely available even in markets where collecting product attributes is challenging. Given the prevalence of unstructured data, the approach offers a valuable addition to the toolbox of empirical researchers. To ease adoption, we provide a Python package, DeepLogit, alongside this paper. [^2] The package is available on PyPI and in the public GitHub repository: github.com/deep-logit-demand/deeplogit. Our package heavily borrows from the existing package Xlogit for GPU-accelerated estimation of mixed logit models, developed by Arteaga et al. (2022).
@@
We validate this approach empirically. A key challenge is that substitution patterns are unobserved, so there is no ground truth to test model predictions. An ideal dataset would let us estimate demand precisely from consumers’ choices and then evaluate how well the model predicts an empirical measure of substitution not used in estimation. To achieve this, we design an experiment in which participants choose a book from a list of options, and we elicit both first and second choices. We randomize prices and rankings of all options. This variation lets us estimate substitution patterns using only first choices and without instruments, while reserving second-choice data for validation. We then test how well the model predicts second choices counterfactually. [^3] We refer to second choice predictions as “counterfactual predictions” in the sense that second choices are not used during estimation, which is based solely on first choices. Because second choices reveal which product a consumer substitutes to if their preferred option is unavailable, predicting them directly validates the model’s ability to recover substitution patterns.
@@
Using these experimental data, we show that both text and images contain useful information about substitution. Intuitively, book covers visually convey information about genre, while descriptions and reviews capture plot details that identify similar books even within the same genre. Our model selection algorithm chooses a review-based mixed logit. This selected model improves counterfactual predictions of second-choices relative to a standard attribute-based mixed logit. In fact, the improvement relative to that model is on par with the improvement that the attribute-based mixed logit achieves over plain logit without random coefficients. Thus, using unstructured data alone, our approach recovers substitution patterns substantially better than standard attribute-based models.
@@
We further validate our approach using observational data from 40 product categories offered on Amazon.com, including groceries, pet food, office supplies, beauty products, electronics, video games, and clothing. We combine online purchases from the Comscore Web Behavior panel with product images, titles, descriptions, and reviews collected from Amazon’s product pages. We show that text and images contain useful signals about substitution in all studied categories, highlighting that our approach is broadly applicable across markets. We find that text performs best in some categories where images might seem ex ante more relevant (e.g., clothing), and vice versa. Thus, it is helpful to collect both text and image data even when only one seems relevant a priori.
@@
Our paper contributes to the extensive literature on demand estimation. We build on papers that model substitution through the heterogeneity of preferences over observed product attributes (Berry et al., 1995; McFadden and Train, 2000; Nevo, 2001; Berry et al., 2004). We show how researchers can incorporate unstructured data into these standard models to improve estimation of substitution patterns. Our key contribution is to validate this approach extensively using both experimental and observational data.
@@
We also build on a vast computer science literature that transforms unstructured data into lower-dimensional representations (Krizhevsky et al., 2012; Mikolov et al., 2013; Goodfellow et al., 2016) and a growing economics and marketing literature that applies pre-trained deep learning models (Battaglia et al., 2024; Dew, 2024). Unlike this prior work, which focuses on predicting observed outcomes, we aim to predict substitution patterns counterfactually—a crucial task for applications like merger simulations and optimal pricing. Because substitution patterns are usually unobserved, we cannot apply standard cross-validation and instead need to design an experiment to measure them. To our knowledge, we are the first to use such an experiment to validate demand models. [^4] Berry et al. (2004) show that second choice data are useful for estimating substitution patterns, and Conlon et al. (2022) estimate demand from aggregate data on first-choice probabilities and a subset of second-choice probabilities. Conceptually, our validation approach reverses the logic of these papers: we only use first-choice data for estimation and reserve second choices for model validation. In a similar spirit to our approach, Raval et al. (2022) use hospital closures induced by natural disasters to compare the ability of various models to predict diversions.
Several other papers leverage text and image data in demand estimation. Quan and Williams (2019) and Lee (2024) incorporate product images and textual descriptions as mean utility shifters. Our approach models utility covariances and can thus capture substitution patterns more flexibly. Han and Lee (2025) use image data to estimate substitution among fonts and study the effects of copyright protection policies. We contribute by validating our approach on ground truth substitution patterns, comparing counterfactual prediction performance across several embedding models and types of unstructured data, and providing practical guidance for demand estimation. Sisodia et al. (2024) extract interpretable product attributes from images and use them to estimate preferences in a choice-based conjoint experiment. Because our approach is easier to implement and scales better across categories, it is preferable when researchers do not need the extracted attributes to be interpretable.
@@
An alternative way to learn substitution patterns is to use survey data. For example, Dotson et al. (2019) ask survey participants to rate each product image and then incorporate rating correlations as shifters of utility correlations into a demand model. Similarly, Magnolfi et al. (2022) elicit relative rankings from survey participants (e.g., “Product A is closer to B than to C”), generate embeddings from these rankings, and use them in a mixed logit model. Our approach complements these survey-based methods but has the advantage of using only widely available text and image data, thereby avoiding the need for costly category-specific surveys.
@@
Our approach is well-suited for empirical applications that require accurate and flexible estimates of product substitution. This includes evaluating how horizontal mergers (Nevo, 2000; Federal Trade Commission, 2022), new product launches (Hausman, 1994; Petrin, 2002), corrective taxes (Allcott et al., 2019; Seiler et al., 2021), and trade restrictions (Goldberg, 1995; Berry et al., 1999) influence consumers’ choices and welfare through changes in assortments and prices. It can also be used to study how multi-product firms should optimize prices and promotions (Hoch et al., 1995; Vilcassim and Chintagunta, 1995). The approach avoids the need to collect category-specific attributes, making it easier to estimate demand across many categories (Döpper et al., 2024). [^5] Döpper et al. (2024) remark that estimating demand across multiple categories “would be difficult to implement at scale because it would require category by-category assessments about which characteristics are appropriate to include and whether or not relevant data are available.” Our approach can address this problem if researchers have access to unstructured data. Finally, this approach might be particularly useful for estimating demand for creative goods—such as books, art, and movies—where unstructured data is likely to capture product differentiation more fully than observed attributes alone.
@@ 2 Proposed Approach
To extract information from product images, we compute low-dimensional representations—embeddings—using pre-trained deep learning models. We use four convolutional neural networks: VGG19 (Simonyan and Zisserman, 2015), ResNet50 (He et al., 2016), InceptionV3 (Szegedy et al., 2016), and Xception (Chollet, 2017). [^6] These models were originally trained to classify images into labeled categories (e.g., “cup,” “book,” or “sofa”). However, since our goal is not to label products but to measure visual features that distinguish them from each other, we remove the classification layer from these models and work directly with embeddings. The key advantage of using pre-trained models is that it reduces the computational burden of estimation while leveraging models that were trained on large-scale datasets and are known for their strong performance in image classification tasks (Russakovsky et al., 2015). Because these models perform well at distinguishing visually similar objects, we expect embeddings to capture the key visual features that differentiate products. Rather than committing to a single model, we perform model selection to determine which one best captures substitution patterns, as described below.
We also extract information from texts, including product titles, descriptions, and reviews. We use a bag-of-words model, which represents text as fixed-length vectors based on word counts, as well as a variation with a TF-IDF vectorizer. [^7] TF-IDF assigns more weight to words that are frequent in a specific text but rare across others, thus emphasizing unique words. This makes it more likely that text embeddings capture distinctive words that differentiate products from one another. While these count models are relatively simple, we view them as useful benchmarks because they can still detect attributes mentioned in textual product descriptions. In addition, we employ two pre-trained deep learning models: Universal Sentence Encoder (USE) (Cer et al., 2018), and BERT Sentence Transformer (ST) (Reimers and Gurevych, 2019). [^8] The model trained by Reimers and Gurevych (2019) is a more efficient version of the widely used BERT network (Devlin et al., 2018). Both USE and ST produce semantically meaningful sentence embeddings and achieve excellent performance on semantic textual similarity benchmarks (Cer et al., 2017). As a result, they can identify when sellers or consumers describe the same product attributes and functional benefits using similar language, even if the exact words differ.
@@
Although embeddings compress texts and images into lower-dimensional representations, they remain high-dimensional compared to the number of attributes typically included in demand models. For example, incorporating 512-dimensional VGG19 embeddings into a random coefficients logit model would require prohibitively costly numerical integration over all dimensions. We apply PCA to further reduce dimensionality (Backus et al., 2021).
PCA is appealing for several reasons. Because embeddings are trained on general-purpose datasets to classify images into broad categories, they contain variation that distinguishes each category from one another (e.g., laptops from tablets). By applying PCA within each category, we downweight this common variation and focus on differences across products within that category. Additionally, the resulting principal components are orthogonal to each other, which avoids multicollinearity issues and thus simplifies estimation of logit demand models with random coefficients.
@@
We include principal components into a standard mixed logit model (Berry et al., 1995, 2004), estimating a separate random coefficient for each included component. Since each deep learning model and data type produces different principal components, a key question is which to include in the demand model. [^9] While we could select the specification that best matches observed second choices in our experiment, this strategy is unavailable to most researchers who usually only have first-choice data. We thus propose a model selection algorithm that relies solely on first choices. We perform model selection using in-sample $AIC$, which balances model fit and complexity and can be interpreted as approximating expected out-of-sample predictive performance (Akaike, 1998). In our experimental results, in-sample $AIC$ strongly correlates with counterfactual performance, justifying its use for model selection (see Section 3.3 for a discussion and comparison to $BIC$ and cross-validation).
Algorithm 1 describes our model selection procedure. We define the candidate set as the variables for which random coefficients may be included. In our applications, this set includes price and the first $P$ principal components. We first estimate the plain logit model without random coefficients ($K=0$) and record its $AIC$. For each $K\geq 1$, we estimate all models with random coefficients on $K$ candidate variables and retain the model with the lowest $AIC$. If this $AIC$ is lower than that of the previously selected model, we update the selected specification and increase $K$; otherwise, we stop and select the previous model. We increase $K$ until $AIC$ no longer improves. We repeat this procedure for each specification—defined by data type and embedding model (e.g., USE Reviews)—and choose the specification with the lowest overall $AIC$.
This algorithm avoids costly combinatorial search over all possible specifications. At the same time, it ensures that we do not rely too heavily on the default ordering of principal components, which reflects the explained variance of embeddings but need not match how predictive they are of preferences. For example, substitution patterns may be driven by the second and fourth principal components. In this case, our approach may select a model with random coefficients on those components, but not on the first or third.
@@
9:     Find subset $R_{K}^{*}\subseteq\mathcal{R}$ that minimizes $AIC$ at $AIC_{K}^{*}$
@@
11:       $\text{BestSpec}^{(m)}\leftarrow R_{K}^{*}$ : update the best specification
12:       $\text{BestAIC}^{(m)}\leftarrow AIC_{K}^{*}$ : update the lowest AICAIC
13:     end if
14:   end while
15: end for
16: Choose the best-fitting specification $m^{*}=\arg\min_{m}\text{BestAIC}^{(m)}$
@@
The key feature of our approach is that it uses text and images as proxies for product attributes that shape substitution. Take tablets, for example: product titles may reveal brand, screen size, and camera resolution (e.g., “Apple iPad 10.2-inch 12MP camera”); seller descriptions may highlight whether the tablet is suitable for drawing or gaming; and consumer reviews may mention that a tablet is durable and child-friendly. Similarly, photos may showcase design features like color and casing style. Our approach extracts this information in an automated way and uses it to estimate substitution patterns.
We interpret the relationship between embeddings and substitution as correlational. In particular, our approach does not assume that consumers directly consider product images or textual descriptions when making choices. Under this interpretation, our approach is well-suited for counterfactuals where embeddings can be held fixed, such as optimal pricing or merger simulations. By contrast, it is less appropriate for settings in which counterfactual changes that may alter product design or positioning and, in turn, the embeddings themselves. In such cases, researchers are required to estimate the causal effect of embeddings on demand. Section 5.2 provides additional discussion.
@@
When some data on observed attributes are available, researchers can incorporate them into our approach. For example, they can expand the set of candidate variables in Algorithm 1 to include the observed attributes themselves or, when the number of attributes is large, their principal components. In our experiment, adding observed attributes does not improve counterfactual predictions, indicating that the information in the attributes is subsumed by the unstructured data. See Sections 3.3 and 5.1 for details.